OpenStudy (anonymous):

solve:-2/2x+27/6x=-7/6

OpenStudy (anonymous):

7/(2 x) = -7/6 Divide both sides by 7/2: 1/x = -1/3 Raise both sides to the power of -1: x = -3

OpenStudy (anonymous):

thank you